# Uni-NaVid: A Video-based Vision-Language-Action Model for Unifying Embodied Navigation Tasks

### v1

@@ Abstract
A practical navigation agent must be capable of handling a wide range of interaction demands, such as following instructions, searching objects, answering questions, tracking people, and more. Existing models for embodied navigation fall short of serving as practical generalists in the real world, as they are often constrained by specific task configurations or pre-defined maps with discretized waypoints. In this work, we present Uni-NaVid, the first video-based vision-language-action (VLA) model designed to unify diverse embodied navigation tasks and enable seamless navigation for mixed long-horizon tasks in unseen real-world environments. Uni-NaVid achieves this by harmonizing the input and output data configurations for all commonly used embodied navigation tasks and thereby integrating all tasks in one model. For training Uni-NaVid, we collect 3.6 million navigation data samples in total from four essential navigation sub-tasks and foster synergy in learning across them. Extensive experiments on comprehensive navigation benchmarks clearly demonstrate the advantages of unification modeling in Uni-NaVid and show it achieves state-of-the-art performance. Additionally, real-world experiments confirm the model’s effectiveness and efficiency, shedding light on its strong generalizability.
@@
00footnotetext: First author (e-mail: zhngjizh@gmail.com). $\dagger$ Corresponding authors (e-mail: zhangzz@galbot.com, hewang@pku.edu.cn).
1 Introduction
Embodied navigation is fundamental for supporting various interaction demands. Recently, instruction-following navigation techniques have advanced significantly with the development of multimodal models. However, existing studies [2, 79, 98, 51, 100, 68, 96, 66, 61, 94] either develop individual models for different tasks, e.g., vision-and-language navigation [37, 39], object goal navigation [11], embodied question answering [18], following [97, 30, 56], etc, or explore task unification in a simplified setting (Tab. 1). They all fall short as full-stack navigation models qualified for practical general-purpose deployment.
@@ 1 Introduction
Table 1: Task and setting comparison. Uni-NaVid is developed to address four embodied navigation tasks, generating action outputs in continuous environments. C.E.: Continuous Environment; D.E.: Discrete Environment.
Developing a full-stack navigation model poses significant challenges, not only in unifying task modeling but also in integrating heterogeneous data in joint use. Initial efforts adopt imitation learning (IL) [72, 78, 55] or reinforcement learning (RL) [90, 82] to learn general navigation skills in simulation environments or limited diverse real-world environments. However, due to the limited rendering quality and diversity of simulators, these approaches often encounter the “sim-to-real” gap and suffer from poor generalization across diverse navigation tasks [22, 5, 33]. Recent studies [100, 98, 51, 50, 67] have attempted to achieve a higher degree of unification on top of pre-trained LLMs, but they simplify the problem to some extent by adopting discretized modeling approaches. They rely on pre-defined graphs for learning decision-making, thus sacrificing the output flexibility and resulting in a low upper bound, especially for real-world deployment.
In this work, we build Uni-NaVid, a Vision-Language-Action (VLA) model for unifying diverse commonly demanded navigation tasks (Tab. 1). It takes video streams and natural language as input, adaptively generates actions following instructions, and automatically responds to questions if raised in an end-to-end manner. We make the first endeavor to address three essential questions, as elaborated below, in pursuit of ultimate real-world application.
How to unify different navigation tasks with minimal problem simplification and maximum utilization of data? We target an up-to-date unification modeling in Uni-NaVid where it takes single-view egocentric video streams and natural language as inputs, and end-to-end generate actions or linguistic responses for instructions or questions adaptively. We follow the natural characteristic of Large Language Models (LLMs) to cross-modally interleave the inputs and outputs of different navigation tasks into a single sequence in the token space to enable joint learning with next-token-prediction paradigm for Uni-NaVid. It incorporates four widely demanded navigation tasks, including Vision-and-Language Navigation (VLN), object-goal navigation, embodied question answering, and human following. The corresponding training data are all collected from simulation environments. Besides, we further incorporate Internet real-world data for Video Question Answering (VQA) [7, 44] and video captioning [16] as auxiliary tasks into this unified model to enhance navigation performance and promote sim-to-real generalization. This is a groundbreaking attempt in this field at unified cross-modal, cross-task, and cross-sim-to-real modeling. It achieves impressive synergy in a broad range and results in a full-stack navigation agent.
How to improve the efficiency of such a unification design to enable a deployment-friendly model? In addition to unification modeling, we also prioritize efficient model design for practical deployment. Specifically, we adopt an online token merging mechanism to compress near historical frames with a relatively low ratio while compressing far historical frames with a relatively high ratio. In this way, Uni-NaVid learn compact representations that maintain not only fine-grained spatial information but also structured temporal information, thus speeding up the model inference by reducing the token number. Besides, Uni-NaVid adopts a foresight prediction to generate actions for a future horizon at once instead of step-by-step. This allows us to use an asynchronous strategy for action inference and execution.
How to verify the synergy between different tasks? We conduct comprehensive ablation studies to verify the synergy over different navigation tasks, as well as between simulation and real-world tasks in both simulation and real-world environments. We find they promote each other with unification modeling in such a broad range and help Uni-NaVid attain clear superiority in all sub-tasks [37, 64, 18, 31] compared to recent works. Besides, we also evaluate Uni-NaVid on long-horizon real-world tasks, which include multiple subtasks, and witness it can complete them in a zero-shot manner as shown in Figure 1.
By exploring these three essential questions, we demonstrate the advantages of unified modeling and joint data utilization across different tasks and in sim-to-real transfer. We hope this will foster future progress in developing an up-to-date full-stack navigation agent for general-purpose real-world applications.
2 Related Works
@@
Generalist Embodied Navigation. Embodied navigation [2, 79, 96] requires agents to navigate in unseen environments based on human instructions. There is extensive literature on embodied navigation; here, we focus on four mainstream tasks that involve both visual information and language instructions: Vision-and-Language Navigation [4, 37, 39], Object Goal Navigation [11], Embodied Question Answering [18], and Human Following [31, 56]. Early efforts [72, 78, 55, 82] towards a generalist-embodied navigation model involved multi-task navigation datasets and directly learning navigation skills, showing initial success in multi-task performance. However, these methods experienced performance drops when deployed in novel environments, especially in real-world settings. In recent years, advanced approaches [98, 51, 29, 50, 101, 67] have leveraged the generalization capabilities of large language models to improve multi-task navigation. These models show promising generalizability across navigation tasks but rely on extensive prompting, which impacts time efficiency. In contrast, our video-based large language model is trained end-to-end for multi-task navigation, offering robust generalization and computational efficiency for tasks like human following.
@@ 3 Method
Overview. As shown in Figure 2, the architecture of Uni-NaVid is designed for general-purpose navigation which is composed of two main components: a vision encoder and a large language model (LLM). First, the video is encoded into a sequence of visual tokens via a vision encoder. The visual tokens are spatially and temporally merged, resulting in a compact sequence of visual tokens. Then they are projected into a space aligned with language tokens, which are referred to as observation tokens. As common, the instructions are also tokenized as a set of tokens, called instruction tokens. Both the observation and instruction tokens are concatenated and passed to the LLM, which infers 4 action tokens representing the next four actions.
3.1 Model of Uni-NaVid
Observation encoding. Given the ego-centric video up to time $T$, denoted by $\mathcal{O}_{T}=\{\mathbf{x}_{1}.\cdots,\mathbf{x}_{T}\}$, we encode the video to a sequence of observation tokens. For each frame $\mathbf{x}_{t}$, we first get its visual tokens $\mathbf{X}_{t}\in\mathbb{R}^{N_{x}\times C}$ with a vision encoder (EVA-CLIP [70] in implementation), where $N_{x}$ is the patch number ($N_{x}$ is set to 256) and $C$ is the embedding dimension. Inspired by vision-language models [46, 44], we also perform a vision-language alignment by projecting the visual feature embedding into the language space, which denotes as observation tokens $\mathbf{E}^{V}_{t}\in\mathbb{R}^{N_{x}\times C}$.
$\mathbf{E}^{V}_{t}=P_{V}(\mathbf{X}_{t}),$ (1)
Where $P_{V}$ represents the cross-modality projector. Although observation tokens provide crucial visual information for the agent to plan its next actions, the linearly increasing number of these tokens $(T\times N_{x})$ leads to progressively longer processing times, making it impractical to use them directly for navigation purposes.
Online visual token merging. To balance the efficiency and effectiveness of visual tokens, we design a token merging strategy to retain navigation-relevant visual information for Uni-NaVid. This strategy builds on the key navigation insight that recent observations are more critical, and that visual information between consecutive frames and neighboring pixels could be redundant.
Drawing inspiration from the Atkinson-Shiffrin memory model [6, 69], we categorize visual tokens into current visual tokens $\mathbf{X}_{\text{curr}}$, short-term visual tokens $\mathbf{X}_{\text{short}}$, and long-term visual tokens $\mathbf{X}_{\text{long}}$. These tokens are grouped based on their timestamps relative to the most recent frame. Current visual tokens capture sufficient visual information for the agent to perceive its immediate surroundings and to plan subsequent paths. Short-term visual tokens maintain temporal continuity, preserving relevant information momentarily to support ongoing planning. Long-term visual tokens facilitate the integration and retrieval of knowledge across the entire navigation history. For each category of visual tokens, we apply a grid pooling operation at different pooling resolutions:
@@ 3.1 Model of Uni-NaVid
where GridPool(·) is a grid pooling operation  [44, 95], squeezing the tokens from $N_{x}$ to $\frac{N_{x}}{\alpha^{2}}$, and $B$ (set to $64$) is the length of the buffer of shorter memory. Here, we adopt the $\alpha_{\text{curr}}=2$, $\alpha_{\text{short}}=8$, $\alpha_{\text{long}}=16$, leads to visual tokens as $\mathbf{X}_{\text{curr}}\in\mathbb{R}^{64\times C}$, $\mathbf{X}_{\text{short}}\in\mathbb{R}^{4\times C}$, $\mathbf{X}_{\text{long}}\in\mathbb{R}^{1\times C}$, respectively. These hyperparameters are determined through empirical experimentation and can be adjusted if memory and computational resources are not primary constraints.
In the navigation process, the agent consistently observes new frames. When a new frame at time $T+1$ arrives, we apply grid pooling to the most recent visual tokens at time $T$ and the oldest short-term visual tokens at time $T-B$, then insert them into the short-term and long-term visual tokens, respectively:
@@
In this way, we can online update the short-term visual tokens and long-term visual tokens with previously generated tokens. Furthermore, we compare the cosine similarity of the $\mathbf{X}_{\text{short}\rightarrow\text{long}}$ with the latest long-term visual tokens $\mathbf{X}_{\text{long}}$ at time $T-B-1$. If the similarity is above a given threshold $\tau$, we merge them based on the number of frames previously merged in the latest long-term visual tokens, $K$.
@@
We insert new long-term visual tokens $\mathbf{X}_{\text{short}\to\text{long}}$ if their similarity falls below $\tau$ (empirically set to $\tau=0.95$ [69]), which indicates they contain relatively distinct visual information. This fusion in temporal preserves the observation history in a highly compact form (in a length of $M\ll T-B-1$), with minimal computational cost for merging these tokens. Only frames at the boundaries of the visual token types require computation through parallel pooling operations, making the process computationally efficient and naturally suited for online deployment in real-world navigation applications. Compared to existing video-based large language models [95, 69, 44], this merging strategy significantly reduces processing time, achieving an average of 0.2 seconds per inference. This improvement becomes increasingly notable when handling longer video sequences. A detailed analysis of time efficiency is provided in the Supplementary Materials.
Action planning. For action planning, we follow the existing work [95] and use special tokens <NAV> to inform the LLM to predict navigation actions. Accordingly, we concatenate the observation token sequence $\mathbf{E}^{V}_{1:T}$ (obtained through Eq. 1 with visual tokens of Eq. 2) with special token <NAV> and instruction tokens and input them into the LLM to infer 4 action tokens $\{\mathbf{E}^{A}_{T},\cdots,\mathbf{E}^{A}_{T+3}\}$:
Input: $\{Long\_term\_frames\}\{Shot\_term\_frames\}\\ \{Current\_frame\}<\!NAV\!>\{Instruction\}$ Output: $<\!Action\_0\!>\!<\!Action\_1\!>\!<\!Action\_2\!>\!\\ \!<\!Action\_3\!>$

### v2

@@ Abstract
Embodied Navigation is a fundamental capability for intelligent robots, requiring robots to follow human commands and move autonomously within physical environments. Despite significant advancements, most existing navigation approaches are tailored to specific navigation tasks, such as instruction following, searching objects, answering questions, tracking people, and more. However, the increasing demands on advanced embodied navigation pose the challenge of designing a practical navigation agent that can incorporate multiple navigation tasks naturally and benefits from the synergy between these tasks. To this end, we present Uni-NaVid, a video-based vision-language-action (VLA) model to unify different paradigms of navigation tasks and improve navigation performance by encouraging the synergy among different navigation sub-tasks. This VLA model can directly take natural language instructions and RGB video streams as inputs and output low-level robotic actions in an end-to-end manner. To efficiently process extensive RGB video streams, we propose an online token merge strategy that spatially and temporally consolidates similar visual information which improves the inference speed to 5 Hz. For training Uni-NaVid, we collect 3.6 million navigation data samples across different navigation tasks. Extensive experiments on diverse navigation benchmarks demonstrate that Uni-NaVid achieves state-of-the-art performance within a unified framework by using only ego-centric RGB video as inputs. Additionally, real-world experiments confirm the model’s effectiveness and efficiency, shedding light on its strong generalizability.
@@
00footnotetext: Author (e-mail: zhngjizh@gmail.com). $\dagger$ Corresponding authors (e-mail: zhangzz@galbot.com, hewang@pku.edu.cn).
I Introduction
Embodied navigation [112, 85] is a critical capability for intelligent robots and has drawn significant attention in the robotics community. For successful embodied navigation, robots must be able to move autonomously within physical environments based on human instructions. However, navigation tasks vary significantly, and most existing studies are designed for specific tasks, e.g., vision-and-language navigation [44, 46], object goal navigation [12], embodied question answering [21, 90], and following [113, 36, 68]. Consequently, most current approaches are developed to address only one type of navigation task, often relying on specialized modules and task-specific datasets. This narrow scope limits their applicability to multi-purpose navigation applications and prevents these methods from leveraging potential synergies across diverse navigation tasks.
Developing a versatile navigation model presents significant challenges, as it requires the unification of navigation task modeling and the integration of heterogeneous data for joint use. Initial efforts adopt imitation learning (IL) [85, 93, 66] or reinforcement learning (RL) [106, 97] to learn general navigation skills in simulation environments or limited diverse real-world environments. However, due to the limited rendering quality and diversity of simulators, these approaches often encounter the “sim-to-real” gap and suffer from poor generalization across diverse navigation tasks [27, 5, 40]. Recent studies [120, 114, 61, 60, 78] have attempted to achieve a higher degree of unification using pre-trained large language models (LLMs). However, due to the low frequency of LLM inference, they simplify the problem to some extent by adopting discretized modeling approaches. They rely on pre-defined graphs for decision-making learning, which sacrifices output flexibility and introduces additional challenges for real-world deployment.
In this work, we propose Uni-NaVid, a video-based Vision-Language-Action (VLA) model for unifying diverse commonly demanded navigation tasks (Tab. I). Uni-NaVid  takes egocentric RGB video streams and natural language instructions as inputs, and directly generates low-level actions for navigation in continuous environments. To achieve multi-task navigation while supporting efficient navigation, Uni-NaVid extend video-based VLM [51] by incoprating two key components: (1) an efficient VLA architecture based on an online token merge mechanism, which enables efficient processing of online-captured video streams for LLM inference; and (2) an extensive collection of 3.6M samples across four widely studied navigation tasks. We provide a detailed elaboration below:
During navigation, the agent is required to process a substantial volume of online captured frames, which results in memory overload and computational latency, particularly in LLM-based approaches [111, 61]. To this end, we propose an online token merging mechanism to compress near historical frames with a relatively low ratio while compressing far historical frames with a relatively high ratio. This merging mechanism operates in an on-the-fly manner, maximizing the reuse of previous navigation history. In this way, Uni-NaVid learn compact representations that maintain not only fine-grained spatial information but also structured temporal information, thus speeding up the model inference by reducing the token number. Besides, Uni-NaVid adopts a foresight prediction to generate actions for a future horizon at once instead of step-by-step. This enables Uni-NaVid to achieve 5Hz inference, facilitating the deployment of a non-blocking navigation robot powered by a VLA model in real-world environments (Please refer to the supplementary video).
We aim to build  Uni-NaVid as a versatile multi-task navigation agent, incorporating four widely demanded navigation tasks: vision-and-language navigation, object-goal navigation, embodied question answering, and human following. These tasks are distinct from each other, with varying task settings and objectives. Specifically, for the human-following task, we construct a new language-guided human-following benchmark for data collection and evaluation. Finally, we collect 3.6M navigation samples based on diverse navigation tasks with different simulation environments. Additionally, inspired by the success of manipulation VLAs [9], we further integrate 2.3M real-world internet data samples for Video Question Answering (VQA) [7, 51] and video captioning [19] as auxiliary tasks. This integration aims to enhance scene understanding and promote sim-to-real generalization.
We conduct extensive experiments on benchmarks across the aforementioned four navigation tasks and compared our method with strong baselines specifically designed for each task. Utilizing only RGB video streams and instructions as inputs, our method demonstrates the superiority of a single VLA model across diverse benchmarks, achieving SOTA or SOTA-comparable performance. Furthermore, comprehensive ablation studies validate the synergistic benefits of learning multiple navigation tasks jointly. Finally, real-world experiments demonstrate that Uni-NaVid achieves non-blocking navigation exhibiting impressive robustness in handling diverse instructions and environments. We believe our work serves merely as a starting point for general-purpose navigation, and we will release the full source code to benefit the community.
II Related Works
@@
TABLE I: Task and setting comparison. Uni-NaVid is developed to address four embodied navigation tasks, generating action outputs in continuous environments. C.E.: Continuous Environment; D.E.: Discrete Environment.
@@
Multi-Task Embodied Navigation. Embodied navigation [2, 94, 112] requires agents to navigate in unseen environments based on human instructions. There is extensive literature on embodied navigation; here, we focus on four mainstream tasks that involve both visual information and language instructions: Vision-and-Language Navigation [4, 44, 46], Object Goal Navigation [12], Embodied Question Answering [21], and Human Following [37, 68, 118, 119]. Early efforts [85, 93, 66, 97] towards a generalist-embodied navigation model involved multi-task navigation datasets and directly learning navigation skills, showing initial success in multi-task performance. However, these methods experienced performance drops when deployed in novel environments, especially in real-world settings. In recent years, advanced approaches [114, 61, 35, 60, 121, 78] have leveraged the generalization capabilities of large language models to improve multi-task navigation. These models show promising generalizability across navigation tasks but rely on extensive prompting, which impacts time efficiency. In contrast, our video-based large language model is trained end-to-end for multi-task navigation, offering robust generalization and computational efficiency for tasks like human following.
Embodied Navigation Datasets. To train and evaluate the performance of a policy for embodied navigation tasks, a large body of datasets and corresponding benchmarks have been proposed [23, 123, 59, 64]. These datasets play a crucial role in the embodied navigation community. Here, we review the datasets most relevant to our methods. For vision-and-language navigation, the most widely used datasets are Room-2-Room (R2R) [4] and Room-cross-Room (RxR) [47], which provide navigation instructions and ground truth trajectories of landmarks. We focus on a variant of R2R and RxR in continuous environments, called VLN-CE [44], which is more practical for real-world applications. For object goal navigation, there are several famous benchmarks such as HM3D [69], MP3D [11], and Aithor [124], which are built on various scene environments and simulators. Here, we leverages the HM3D dataset on Habitat [76], which shares the same action settings as VLN-CE. For embodied question answering (EQA), there are diverse datasets focusing on different attributes of EQA, such as MP3D-EQA [90], MT-EQA [103], Graph-EQA [83], and MX-EQA [38]. We select MP3D-EQA, which is well-maintained with the latest baselines. For human-following [116, 117] benchmarks, there is currently no benchmark that provides textual descriptions of humans. Therefore, we have self-built a textual description-based human-following benchmark using Habitat 3.0 [68]. Note that new benchmarks are consistently being proposed, covering a diverse range of navigation attributes. However, our goal is to train and evaluate our method on mainstream datasets to clearly justify the performance of our approach.
@@ III Problem Formulation
Overview. As illustrated in Figure 2, Uni-NaVid  is composed of three main components: a vision encoder, an online token merge mechanism and a large language model (LLM). First, the online captured video stream is encoded by the vision encoder (EVA-CLIP [82] in implementation) to extract frame-wise visual features in the form of tokens, which we denote them as visual tokens. The visual tokens are then spatially and temporally merged by leveraging an online token merge mechanism. Next, the merged visual tokens are projected with an MLP projector into a feature space aligned with language tokens, which are referred to as visual observation tokens. As common, the instructions are also tokenized as a set of tokens, known as language observation tokens. Both the visual observation tokens and language observation tokens are concatenated and passed to the Large Language Model (LLM), which infers four action tokens that represent the next four actions.
IV Model of Uni-NaVid
IV-A Observation Encoding.
Given the ego-centric video up to time $T$, denoted by $\mathcal{O}_{T}=\{\mathbf{x}_{1}.\cdots,\mathbf{x}_{T}\}$, we encode the video to a sequence of visual features in the form of tokens. For each frame $\mathbf{x}_{t}$, we first get its visual feature tokens $\mathbf{X}_{t}\in\mathbb{R}^{N_{x}\times C}$ with a vision encoder (EVA-CLIP [82] in implementation), where $N_{x}$ is the patch number ($N_{x}$ is set to 256) and $C$ is the embedding dimension.
$\mathbf{X}_{1:T}=Encoder(\mathbf{x}_{1:T})$ (1)
The visual features provide rich information that enables the agent to understand its navigation history and plan subsequent actions. However, during navigation, the progressively increasing number of visual tokens $(T\times N_{x})$ results in progressively longer inference times for the LLM (typically 1–2 seconds per inference) [111]. This increased latency renders LLM-based navigation impractical for deployment in real-world environments.
IV-B Online Visual Token Merging
To reduce the number of visual tokens while preserving sufficient navigation visual information, we design an token merging mechanism. This strategy is based on the key insight that recent observations are more critical for navigation, and that visual information between consecutive frames (temporally) and within neighboring pixels (spatially) may be redundant.
Visual token grouping. Drawing inspiration from the Atkinson-Shiffrin memory model [6, 80], we categorize visual tokens into current visual tokens $\mathbf{X}_{\text{curr}}$, short-term visual tokens $\mathbf{X}_{\text{short}}$, and long-term visual tokens $\mathbf{X}_{\text{long}}$. These visual tokens are grouped based on their timestamps relative to the current frame $T$ and for each group of visual tokens, we apply a grid pooling operation at different pooling resolutions:
@@ IV-B Online Visual Token Merging
where GridPool(·) is a grid pooling operation  [51, 111], spatially squeezing the tokens from $N_{x}$ to $\frac{N_{x}}{\alpha^{2}}$, and $B$ (set to $64$) is the length of the buffer of shorter memory. Here, we adopt the $\alpha_{\text{curr}}=2$, $\alpha_{\text{short}}=8$, $\alpha_{\text{long}}=16$, leads to visual tokens as $\mathbf{X}_{\text{curr}}\in\mathbb{R}^{64\times C}$, $\mathbf{X}_{\text{short}}\in\mathbb{R}^{4\times C}$, $\mathbf{X}_{\text{long}}\in\mathbb{R}^{1\times C}$, respectively. Here, current visual tokens $\mathbf{X}_{\text{curr}}$ encapsulate comprehensive visual information, enabling the agent to perceive its immediate environment and plan subsequent trajectories. Meanwhile, $\mathbf{X}_{\text{short}}$ and $\mathbf{X}_{\text{long}}$ capture temporally rich information from the captured video stream, facilitating the agent’s comprehension of its navigation history.
It should be noted that these hyperparameters are obtained through empirical experimentation to achieve an optimal balance between manageable token numbers and adequate visual information representation. These hyperparameters can be further adjusted when memory capacity and computational resources are not limiting factors. We provide a detailed explanation and ablation study of $\alpha$ in the supplemental material.
Online visual token process. During the navigation process, the agent consistently observes new frames. However, performing encoding and grouping (Eq. 2) for all frames at each step would be computationally intensive. To address this, we implement an online visual token processing mechanism that maximizes the reuse of previously generated visual tokens. Specifically, when a new frame at time $T+1$ is received, we apply grid pooling exclusively to the most recent visual tokens at time $T$ and the oldest short-term visual tokens at time $T-B$. These processed tokens are then integrated into the short-term and long-term visual tokens, respectively:
@@
To prevent the linear growth of long-term visual tokens $\mathbf{X}_{\text{Long}}$, we further perform token merging on the long-term visual tokens by combining adjacent tokens that exhibit high similarity, following the approach of VLM-based methods [8, 80]. Specifically, we merge the long-term visual tokens based on the cosine similarity between $\mathbf{X}_{\text{short}\rightarrow\text{long}}$ and the most recent long-term visual tokens $\mathbf{X}_{\text{long}}$ at time $T-B-1$. If the similarity exceeds a predefined threshold $\tau$, we merge them according to the number of frames previously merged (denoted as $K$) in the latest long-term visual tokens:
@@
We insert new long-term visual tokens $\mathbf{X}_{\text{short}\to\text{long}}$ when their similarity falls below a threshold $\tau$ (empirically set to $\tau=0.95$ [80]), indicating that they contain relatively distinct visual information. This online visual token processing preserves the navigation visual history in a highly compact form (with a length of $M\ll T-B-1$). Notably, only visual tokens at the boundaries of groups require parallelizable grid pooling, making the process computationally efficient and naturally suited for online deployment in real-world navigation tasks. We give a description of our token merging technique at Algorithmn 1.
Algorithm 1 Online Visual Token Merging
1:
2:
  • Total number of frames $T$
  • Short memory buffer length $B$
  • Grid pooling scales: $\alpha_{\text{curr}}$, $\alpha_{\text{short}}$, $\alpha_{\text{long}}$
  • Current visual tokens: $\mathbf{X}_{T}\in\mathbb{R}^{N_{x}\times C}$
  • Previously merged tokens: $\mathbf{X}_{\text{curr}}$, $\mathbf{X}_{\text{short}}$, $\mathbf{X}_{\text{long}}$
  • Number of frames merged in the last tokens of long memory: $K$
3:
4:
  • Updated merged tokens: $\mathbf{X}^{\prime}_{\text{curr}}$, $\mathbf{X}^{\prime}_{\text{short}}$, $\mathbf{X}^{\prime}_{\text{long}}$
  • Updated number of frames merged in the last tokens of long memory: $K^{\prime}$
5:
6:if $T==1$ then $\triangleright$ First frame, empty history tokens
7:     $\mathbf{X}^{\prime}_{\text{short}},\mathbf{X}^{\prime}_{\text{long}}\leftarrow[]$
8$\triangleright$ Update short-term visual tokens
9:     $\mathbf{X}_{\text{curr}\rightarrow\text{short}}\leftarrow GridPool(\mathbf{X}_{\text{curr}},\frac{\alpha_{\text{short}}}{\alpha_{\text{curr}}})$
10:     $\mathbf{X}^{\prime}_{\text{short}}\leftarrow\mathbf{X}_{\text{short}}+[\mathbf{X}_{\text{curr}\rightarrow\text{short}}]$
11:end if
12:$\mathbf{X}^{\prime}_{\text{curr}}\leftarrow GridPool(\mathbf{X}_{T},\alpha_{\text{curr}})$triangleright New current visual token
13:if $T>B+1$ then $\triangleright$ Out of short-term tokens buffer
14:     $\mathbf{X}_{\text{short}\rightarrow\text{long}}\leftarrow GridPool(\mathbf{X}_{\text{short}}[0],\frac{\alpha_{\text{long}}}{\alpha_{\text{short}}})$
15:     $\mathbf{X}^{\prime}_{\text{short}}\leftarrow\mathbf{X}_{\text{short}}[1:]$
16:     $s\leftarrow\cos(\mathbf{X}_{\text{long}}[-1],\mathbf{X}_{\text{short}\rightarrow\text{long}})$
17:     if $T>B+2\ \text{and}\ s>\tau$ then $\triangleright$ Fuse long-term tokens
18:         $\mathbf{X}_{\text{last\_long}}\leftarrow\frac{1}{K+1}(K\mathbf{X}_{\text{long}}[-1]+\mathbf{X}_{\text{short}\rightarrow\text{long}})$
19:         $\mathbf{X}^{\prime}_{\text{long}}\leftarrow\mathbf{X}_{\text{long}}[:-1]+[\mathbf{X}_{\text{last\_long}}]$
20:         $K^{\prime}\leftarrow K+1$
21:     $\triangleright$ Add new long-term token
22:         $\mathbf{X}^{\prime}_{\text{long}}\leftarrow\mathbf{X}_{\text{long}}+[\mathbf{X}_{\text{short}\rightarrow\text{long}}]$
23:         $K^{\prime}\leftarrow 1$
24:     end if
25:end if
Compared to existing video-based large language models [111, 80, 51], this online merging strategy significantly reduces inference time, achieving an average of 0.2 seconds per inference. This improvement becomes increasingly notable when handling longer video sequences. A detailed analysis of time efficiency is provided in the Supplementary Materials.
IV-C Action Planning
After obtaining the merged visual tokens from semantic features  [82], we adopt established practices in Vision-and-Language models  [54, 51] to perform vision-language alignment, enabling the large language model (LLM) to effectively interpret visual information. Specifically, we leverage a cross-modality projector $P_{V}(\cdot)$ to project all merged visual tokens $X_{\text{merged}}=\{\mathbf{X}_{\text{long}},\mathbf{X}_{\text{short}},\mathbf{X}_{\text{curr}}\}$ into visual observation tokens that are compatible with the LLM’s input representation space:
$\mathbf{E}^{V}_{T}=P_{V}(\mathbf{X}_{\text{merged}}),$ (7)